# NL-SLAM for OC-VLN: Natural Language Grounded SLAM for Object-Centric VLN

### v1

@@ IV-C Data association of inferred landmarks and observations
where $\mathcal{F}_{t}(.)$ provides text features by encoding the landmark name in the inferred graph, and $\mathcal{F}_{o}(.)$ gives text features of the observation. If the similarity is below a threshold no association is made. This approach offers a balance between effectiveness and computational efficiency. Future iterations of the system may explore more sophisticated data association techniques to handle ambiguous cases.
@@ IV-D Navigation Policy
Localizing the robot within the language-inferred map enables a straightforward navigation policy: The robot navigates by sequentially moving to inferred waypoints, transitioning when within 0.5m of each, and stopping at the last one. However, high uncertainty in waypoint locations can lead to inaccurate navigation. In these cases, the robot should switch to exploration mode, seeking landmarks to reduce uncertainty about upcoming waypoints.
To balance exploration and exploitation, we employ the following strategy for waypoint selection:
@@
This equation selects the inferred waypoint $w_{j}$ that minimizes two factors: the distance between the waypoint and the current robot pose ($||w_{j}\ominus x_{i}||^{2}$), and the uncertainty of the waypoint, represented by the trace of its information matrix $\Lambda=\Sigma^{-1}$. The weighting constant $\alpha$ balances the desire to move towards nearby waypoints with the need to reduce uncertainty.
To determine the precise navigation target, we further sample a point from the posterior distribution $p(w^{*})$ within the navigable region around $w^{*}$. Initially, the posterior distributions of waypoints have high uncertainty, promoting exploration. As the robot gathers observations, uncertainty decreases, refining navigation targets. This adaptive approach allows the robot to navigate efficiently while continuously improving its spatial understanding.
@@ V Experiments
Metrics. We use the commonly used evaluation metrics for VLN [21, 77, 78], such as Success Rate (SR), Success weighted by inverse Path Length (SPL), Oracle Success Rate (OSR), and normalized Dynamic-Time Warping (nDTW). The first three metrics measure the robot’s performance based on how well it can reach the target object, while nDTW measures how closely it can follow the instructions.
Implementation. We evaluate NL-SLAM on OC-VLN in the Habitat simulator [70] in a zero-shot manner. We use a pipeline of pre-trained open-vocabulary object detection for landmark recognition. Given an RGB image, we use RAM [79] for identifying what objects are in the image, Grounding-Dino [19] to predict bounding boxes, and finally, Segment-Anything (SAM) [20] to obtain the semantic segmentation. For the navigator, we use a PointNav policy [80] pre-trained on HM3D [81] scenes. We compare our method to two baselines, both SOTA zero-shot navigation methods for related tasks that have been successful in the real world.
Seq-VLFM baseline. VLFM [82] is an Object-Goal navigation method, which we execute sequentially. We use GPT-4 [15] to extract the objects from the instruction and then run VLFM to go to each object in order. Using BLIP-2 [83], VLFM creates value maps of likely object locations, guiding the frontier-based exploration [84] to efficiently find objects.
InstructNav baseline. InstructNav [85] is a method for general navigation instruction following which evaluates on Object-Goal navigation, VLN, and demand-driven navigation tasks. InstructNav uses an LLM to break the navigation instruction down into a series of paired actions and landmarks. It then plans how to reach each landmark using a combination of value maps, which represent key information about the environment and its relation to the robot, such as semantic information, actions, trajectory history, and intuition from the LLM. They replan with the LLM at each decision-making step to better fit the plan to the observed environment.
@@ VI Results
Benchmark results. Table II shows that NL-SLAM outperforms the baseline methods on all metrics. Our performance on the nDTW metric (48%) specifically indicates that our method is able to follow instructions better than the others. This is intuitive since our method transforms the instructions into a graph, which is then used to perform the navigation, compared to Seq-VLFM, which searches for objects by extracting them from the instructions and thus disregarding the actions and relations encoded into them. Being able to follow the instructions also enables NL-SLAM to successfully reach the target object by following the suggested path as indicated by its performance on the SR and SPL metrics. The OSR metric reflects whether the agent has been near the target object and failed to stop. We find that all the methods achieve reasonable OSR on the dataset. However, the gap between SR and OSR is smaller for NL-SLAM (20%) compared to the others (42% in Seq-VLFM and 43% in InstructNav), indicating that our method is able to call the Stop action correctly more often compared to the others.
TABLE II: Performance. NL-SLAM significantly outperforms Seq-VLFM and InstructNav on all metrics.
@@
NL-SLAM (Ours) 0.58 0.40 0.78 0.48
Qualitative analysis. In Fig. 5, we visualize our agent performance on a single episode over time. At the beginning of the episode ($t=1$), when the agent has not yet made observations in the world, the landmark nodes ($l_{i}=\{l_{1}=``brown~{}couch",l_{2}=``fireplace"\}$) and the waypoint nodes ($w_{i}=\{w_{2},w_{3}\}$) have large uncertainties (visualized as ellipses). As the episode progresses, the agent makes the first observation at $t=20$ and grounds $l_{1}$ and $w_{1}$, thereby becoming certain about their locations. At $t=50$, the agent observes $l_{2}$, thus reducing uncertainty for all the nodes in the inferred graph. Finally, the agent generates the Stop action when it is within 0.5m of the last waypoint $w_{2}$ in the graph.
Figure 5: NL-SLAM in action. Visualizing the progress of our agent through an episode shows how the language-inferred graph gets optimized over time $t$, leading to a successful completion. (left) shows detected objects; (right) shows current robot pose, ground-truth trajectory (green), inferred waypoints (yellow) and landmarks (blue).
To compare NL-SLAM and Seq-VLFM, we visualize the agent trajectories for different episodes (Fig. 6) and observe that the agent follows the instruction better in NL-SLAM compared to Seq-VLFM. This is evident from how closely the agent path matches the ground-truth path in our method.
@@
Ablations. We perform ablations on the different modules in NL-SLAM (Table III) to understand the contribution of each. For both Object Detector and Navigator, we find using an Oracle performs the best, as is expected. The Oracle detector+navigator achieves a +12% increase in success, +13% increase in SPL, and +10% increase in nDTW, indicating that the performance of our method is limited by the choice of the detector and the navigator and can be improved further.
TABLE III: Ablations. Oracle versions of the Object Detector and the Navigator perform better than the pre-trained models, indicating the scope of improvement for NL-SLAM.
@@
Failure analysis. We note two main failure cases: (a) the presence of multiple instances of the same object category; and (b) the misclassification of an object. In the first case, an instruction ‘move through an open door’ while there are multiple doors (Fig. 7) makes it difficult for the agent to figure out the correct ‘door’. The second case is a limitation of the object detector where it incorrectly identifies a ‘television’ as ‘fireplace’, leading to erroneous data association.
Figure 7: Failure cases. Two frequent failure cases in NL-SLAM are: (top) multiple instances of the same landmark present, e.g. ‘door’; and (bottom) an object is mis-classified e.g. ‘television’ identified as ‘fireplace’.
Real-world demonstrations. We deployed NL-SLAM on a Spot robot from Boston Dynamics to demonstrate its effectiveness in the real world. We replaced the PointNav policy with the Boston Dynamics Spot SDK as the navigator to move the robot to a waypoint. We use RTAB-Map to obtain the robot’s coordinates and the observations in a reference frame. Instead of our object detection pipeline, we use YOLOv7 [86] for faster real-time execution. We ran NL-SLAM in an office space with commonly occurring objects such as chairs, tables, potted plants, etc., and demonstrated that our method was able to perform well in this task.
@@ VII Conclusion
We present OC-VLN, a dataset for object-centric vision-and-language navigation enabling us to study the distinct challenge of object-based navigation instruction following. To address this task, we propose NL-SLAM, which robustly grounds objects in the instruction to observations in the environment using a pose graph. NL-SLAM creates an inferred pose graph from the instruction using an LLM, thus incorporating relevant information as a prior for the pose graph. A comparison to two baselines, Seq-VLFM [82] and InstructNav [85], which are SOTA on related tasks, shows that NL-SLAM outperforms them on all metrics.
NL-SLAM can struggle to identify the relevant objects when multiple of the same class are present. This could be addressed by incorporating scene layout priors to predict likely locations for landmarks that have not yet been observed and thus identifying promising directions for the path as a whole. Actively selecting poses to obtain multiple good views of the object, perhaps drawing inspiration from active SLAM, is an avenue to address the other significant cause of failure identified for NL-SLAM, i.e. object misclassification. More sophisticated data association techniques may also alleviate these limitations.
While object-centric vision-and-language navigation is far from solved as a distinct problem, future work should also investigate modularly including more aspects present in VLN. For example, regions as landmarks as well as objects.

### v3

@@ IV-C Data association of inferred landmarks and observations
where $\mathcal{F}_{t}(.)$ denotes the landmark features in the inferred graph, and $\mathcal{F}_{o}(.)$ denotes the observation features. If the similarity is below a threshold no association is made. This simple approach is effective and computationally efficient, future work may explore more sophisticated data association techniques to resolve ambiguity better.
@@ IV-D Navigation policy
Localizing the robot within the language-inferred graph enables us to use a straightforward navigation policy. The robot navigates by sequentially moving to inferred waypoint nodes in the graph, transitioning when within 0.5 meter of each, and stopping at the last node. However, high uncertainty in waypoint locations can lead to inaccurate navigation. In these cases, the robot should switch to exploration mode, seeking landmarks to reduce uncertainty about upcoming waypoints. To balance exploration and exploitation, we employ the following strategy for waypoint selection:
@@
This equation selects the inferred waypoint $w_{j}$ that minimizes two factors: the distance between the waypoint and the current robot pose ($||w_{j}\ominus x_{i}||^{2}$), and the uncertainty of the waypoint, represented by the trace of its information matrix $\Lambda=\Sigma^{-1}$. The balance between these factors is controlled by the constant $\alpha$.
To determine the precise navigation target, we further sample a point from the posterior distribution $p(w^{*})$ within the navigable region around $w^{*}$. Initially, the posterior distributions of waypoints have high uncertainty, promoting exploration. As the robot gathers observations, uncertainty decreases, thus refining navigation targets.
@@ V Experiments
Task. At each time step, the robot has access to RGB and depth images, as well as the entire language instruction and is expected to output one of four actions: move forward, turn left, turn right and stop. We use the same agent embodiment and task setting as VLN-CE [18].
Metrics. We use standard evaluation metrics for VLN [16, 75, 76]: Success Rate (SR), Success weighted by inverse Path Length (SPL), Oracle Success Rate (OSR), and normalized Dynamic-Time Warping (nDTW). The first three measure whether the robot can reach the target object, while nDTW measures how closely it can follow the instructions.
Implementation. We evaluate LIFGIF on OC-VLN in the Habitat simulator [65] in a zero-shot manner. We use a pipeline of pre-trained methods for landmark recognition. We use RAM [77] to identify objects, followed by Grounding-Dino [14] to predict bounding boxes, and finally, Segment-Anything (SAM) [15] to obtain the semantic segmentation. For the navigator, we use a PointNav policy [40] pre-trained on HM3D [78] scenes. We compare to two zero-shot baselines, a SOTA ObjectNav method adapted to our task and an LLM-based SOTA method for instruction-following:
Seq-VLFM baseline. VLFM [48] is a single-object navigation method, which we execute sequentially for all objects, extracted using GPT-4 [11], from instructions. Using BLIP-2 [79], VLFM creates value maps of likely object locations, guiding the frontier-based exploration [80] to efficiently find objects. Note that this baseline completely disregards fine-grained navigation instructions on how to reach the target.
InstructNav baseline. InstructNav [81] is a method for general navigation instruction following which evaluates on Object-Goal navigation, VLN, and demand-driven navigation tasks. InstructNav uses an LLM to break navigation instructions into a series of paired actions and landmarks. It then plans with an LLM on how to reach each landmark using a combination of value maps representing key information about the environment such as semantic information, actions, trajectory history, and intuition. While InstructNav uses LLM as a planner, our method efficiently uses LLM offline to generate inferred graphs from instructions just once.
@@ VI Results
Benchmark results. Table II shows that LIFGIF outperforms the baseline methods on all metrics. Our performance on the nDTW metric (48%) specifically indicates that our method is able to follow instructions better than the others. This is intuitive since our method transforms the instructions into a graph, which is then used to perform the navigation, compared to Seq-VLFM, which searches for objects mentioned in the instructions by disregarding the actions and relations encoded into them, and InstructNav, which uses LLM to directly plan over the instructions. Being able to follow the instructions also enables LIFGIF to successfully reach the target object as indicated by its performance on the SR and SPL metrics. The OSR metric reflects whether the agent has been near the target object but failed to stop. We find that all the methods achieve reasonable OSR, indicating that the objects are easy to navigate to. However, the gap between SR and OSR is smaller for LIFGIF (20%) compared to the others (42% in Seq-VLFM and 43% in InstructNav), indicating that our method is able to call the Stop action correctly more often. This shows that the simple and straightforward stopping criteria in LIFGIF works fairly well, although a more sophisticated stopping policy may yield improvement.
TABLE II: Performance. LIFGIF significantly outperforms Seq-VLFM and InstructNav on all metrics.
@@
LIFGIF (Ours) 0.58 0.40 0.78 0.48
Qualitative analysis. In Fig. 4, we visualize our agent performance on a single episode over time. At the beginning of the episode ($t=1$), when the agent has not yet made observations in the world, the landmark nodes, $l_{i}=\{l_{1}=``brown~{}couch",l_{2}=``fireplace"\}$ and the waypoint nodes, $w_{i}=\{w_{2},w_{3}\}$ have large uncertainties (visualized as ellipses). As the episode progresses, the agent makes observation and grounds $l_{1}$ and $w_{1}$, thereby becoming certain about their locations. At $t=50$, the agent observes $l_{2}$, thus reducing uncertainty for all the nodes in the inferred graph. Finally, the agent generates the Stop action when it is within 0.5m of the last waypoint $w_{2}$ in the graph.
Figure 4: LIFGIF in action. Visualizing the progress of our agent through an episode shows how the language-inferred graph gets optimized over time $t$ by making observations in the real-world and performing data association, leading to a successful completion. (left) shows detected objects; (right) shows current robot pose, ground-truth trajectory (green), inferred waypoints (yellow) and landmarks (blue).
To compare LIFGIF and Seq-VLFM, we visualize the agent trajectories for different episodes (Fig. 5) and observe that LIFGIF follows the instruction better compared to Seq-VLFM. This is evident from how closely the agent path matches the ground-truth path in our method.
@@
Ablations. We perform ablations on the different modules in LIFGIF (Table III) to understand the contribution of each. For both Object Detector and Navigator, we find using an Oracle performs the best, as is expected. The Oracle detector+navigator achieves a +12% increase in success, +13% increase in SPL, and +10% increase in nDTW, indicating that the performance of our method is limited by the choice of the detector and the navigator and can be improved further.
TABLE III: Ablations. Oracle versions of the Object Detector and the Navigator perform better than the pre-trained models, indicating the scope of improvement for LIFGIF.
@@
Failure analysis. We note two main failure cases: (a) the presence of multiple instances of the same object category; and (b) the misclassification of an object. In the first case, an instruction ‘move through an open door’ in the presence of multiple doors (Fig. 6) makes it difficult for the agent to figure out the correct ‘door’. The second case is a limitation of the object detector where it incorrectly identifies a ‘television’ as ‘fireplace’, leading to erroneous data association.
Figure 6: Failure cases. Two frequent failure cases in LIFGIF are: (top) multiple instances of the same landmark present, e.g. ‘door’; and (bottom) an object is mis-classified e.g. ‘television’ identified as ‘fireplace’.
Real-world demonstrations. We deployed LIFGIF on a Spot robot from Boston Dynamics to demonstrate its effectiveness in the real world. We replaced the PointNav policy with the Boston Dynamics Spot SDK as the navigator to move the robot to a waypoint. We use RTAB-Map to obtain the robot’s coordinates and the observations in a reference frame. Instead of our object detection pipeline, we use YOLOv7 [82] for faster real-time execution. We ran LIFGIF in an office space with commonly occurring objects such as chairs, tables, potted plants, etc., and demonstrated that our method was able to perform well in this task.
@@ VII Conclusion
We introduce the OC-VLN dataset to study the distinct challenge of object-centric instruction following. Additionally, we propose LIFGIF, which robustly grounds dense language instructions in unseen visual environments, while navigating in a zero-shot manner. We achieve this via a combination of LLM-based technique and traditional navigation techniques from robotics. Moreover, compared to prior works that use LLM as a planner, we use it much more efficiently to generate inferred graph from language instructions offline and use the widely used factor-graph based SLAM to gradually optimize the graph during navigation. We demonstrate through experiments that LIFGIF outperforms two zero-shot state-of-the-art baselines across all metrics.
We find that LIFGIF can struggle to identify the relevant object in the presence of multiple objects of the same class. Future work in this direction could explore backtracking strategies to recover from such failures. Actively selecting poses to obtain multiple good views of the object, perhaps drawing inspiration from active SLAM, could be another avenue for future investigation to address the other significant cause of failure in LIFGIF, i.e. object misclassification.
While object-centric vision-and-language navigation is far from solved as a distinct problem, future work in this area should also explore including other distinctive environmental aspects such as regions as landmarks in the instructions, making the task more challenging.